OpenStudy (anonymous):

x cos(4x)dx u = x dv = cos 4xdx @LeibyStrauss

OpenStudy (welshfella):

use integration by parts u = dx ---> du = 1 dv = -cos4xdx --- > v = - (1/4) sin 4 x INT udv = uv - INT vdu

OpenStudy (anonymous):

\[\int\limits_{}^{}\cos(4x)dx\] u = 4x du = 4dx 1/4du = dx \[\frac{ 1 }{ 4 }\int\limits_{}^{} \cos(u)dx = \frac{ 1 }{ 4 }\sin(4x)\] \[v = \frac{ 1 }{ 4 }\sin(4x)\]

OpenStudy (anonymous):

\[\int\limits_{}^{} u dv = uv- \int\limits_{}^{} v du\]

OpenStudy (anonymous):

\[x \frac{ 1 }{ 4 } \sin(4x) - \int\limits \frac{ 1 }{ 4 } \sin(4x)\] now let's solve \[\int\limits \frac{ 1 }{ 4 } \sin(4x)\] u = 4x du = 4dx 1/4 du = dx \[\frac{ 1 }{ 4 }\int\limits \sin(u)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }[-\cos(u)] = -\frac{ 1 }{ 4 }\cos(4x)\]

OpenStudy (danjs):

reverse of the product rule on the derivatives [f*g] ' = f ' *g +g ' *f parts is the integration of that

OpenStudy (anonymous):

@vkarwee I made a mistake somewhere.

OpenStudy (anonymous):

i think you forgot to multiply

OpenStudy (anonymous):

I forgot to pull out the 1/4 and then multiply by 1/4 again. I'll redo it

OpenStudy (anonymous):

\[1/4\int\limits_{}^{}\sin(4x)\]

OpenStudy (danjs):

sin(4x) / 16 + x* cos(4x)/4

OpenStudy (anonymous):

= \[= -\frac{ 1 }{ 16 }\sin(4x)\]

OpenStudy (danjs):

sometimes, if you can have the answer, with a calculator or whatever first, it may help you figure out the method, from the answer form

OpenStudy (anonymous):

Final answer \[\frac{ x }{ 4 } \sin(4x) + \frac{ 1 }{ 16 }\cos(4x)\]

OpenStudy (danjs):

the sin and cos is mixed around,

OpenStudy (anonymous):

@DanJS v = 1/4 sin(4x) \[\int\limits v du = \frac{ 1 }{ 16 } \cos(4x)\]

OpenStudy (danjs):

forgot a the minus

OpenStudy (danjs):

forget i was evfen at this question, i did the original prob x*cos(4x) to get that answer, intresting how they are related though

OpenStudy (danjs):

**x*sin(4x) is what i started with my mistake

OpenStudy (danjs):

if you need to know integral of x*sin(4x) , you know how it is related to the cos version